y=f(x)is a linear equation. If f(1)=2 and f(3)=7, find the slope.
Answer by jim_thompson5910(35256)

m = (f(x1) - f(x2))/(x1 - x2)

m = (f(1) - f(3))/(1 - 3)

m = (2 - 7)/(1 - 3)

m = (-5)/(-2)

m = 5/2

So the slope is 5/2